OpenStudy (anonymous):

While docked, the captain wants to resupply the ship’s missiles and ray-guns. You will need to explain to the captain what information you must know and how that can be used to determine the number of missiles and ray-guns to fill the ship and use all the batteries. The ship can hold only 15 weapons total and only has 45 batteries. The missiles and ray-guns each require different amounts of batteries to operate.

OpenStudy (anonymous):

This is what my teacher said because i got it wrong: 4) This one is asking what else you need to know. What do you need to know about the batteries? From there you can set up a system of equations. Where is the solution?

OpenStudy (anonymous):

@IMStuck

OpenStudy (imstuck):

Ok, let me look at it for a minute, ok? You're supposed to do what exactly? Write an equation or something? Tell me while I look at it.

OpenStudy (anonymous):

Yes a system of equations and solve for how many batteries a ray gun and a missile needs.

OpenStudy (imstuck):

Ok first you will relate the number of missiles and ray guns to the total number of weapons the ship can hold, which is 15 total. THEN you need to relate the number of batteries used in each type of weapon to the total number of batteries on the ship, which is 45. Let's call missiles "m" and ray guns "r". The first equation is this: m + r = 15. That says that the number of missiles plus the number of ray guns equals 15.

OpenStudy (imstuck):

See that?

OpenStudy (anonymous):

Ok yes I see it.

OpenStudy (imstuck):

Now let me work on the other equation for a sec before I open my big mouth about it, ok? ; )

OpenStudy (anonymous):

Ok

OpenStudy (imstuck):

Solving it to see that the whole system works...BRB

OpenStudy (anonymous):

Ok

OpenStudy (imstuck):

From the problem the way it is stated, which is a bit unclear, to tell you the truth, I am getting a second equation of: mx + rx = 45 That says that x is the number of batteries required per missile and the number of batteries required per ray gun equals 45. But do the weapons both use the same type of battery? Just a different number?

OpenStudy (anonymous):

but the missiles and ray-guns require a different amount of batteries

OpenStudy (imstuck):

Here's what I did and my interpretation of the scenario: m + r = 15 mx + rx = 45 r = 15 - m so... mx + (15 - m)x = 45 mx + 15x - mx = 45 15x = 45 x = 3 Now I am not sure if that is correct (we might need three equations because of the 3 unknowns!)

OpenStudy (imstuck):

Let me think on this for a second...please?

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

The whole idea of this problem is that you are not supposed to actually solve it! Here's the proof from your problem above: "You will need to explain to the captain what information you must know and how that can be used to determine the number of missiles and ray-guns to fill the ship and use all the batteries." that was copied and pasted from the problem. The thing you need to know in order to solve this problem is the number of batteries each weapon requires. You can set up a system of equation with 3 unknowns, but you won't be able to solve it til the captain knows how many batteries each weapon requires. That's your x, the number of batteries in each weapon.

OpenStudy (imstuck):

System is still the same system: m + r = 15 (that's apparent and easy to solve for either m or r) mx + rx = 45 (this is the one that is the trouble. We don't know how many batteries each weapon takes, so we cannot solve the system.) The answer is that the captain needs to tell you the number of batteries each weapon requires, ok?

OpenStudy (anonymous):

ok thankyou